OpenStudy (yttrium):

Hey. Anybody wants to check this out. Solve the initial value problem. y'' + 4y' + 13y = 2e^(-t)

OpenStudy (yttrium):

Condition is y(0) = 0 and y'(0) = 1

OpenStudy (shamil98):

fun such fun. write out the auxillary equation, while i pull up my trustie calc book..

OpenStudy (yttrium):

What I did is: [s^2 F(s) - s y(0) - y'(0) + 4 [ s F(s) - y(0)] +13 F(s) = 2/(s+1) (s^2 + 4s + 13) F(s) = 3+s/(s+1)

OpenStudy (yttrium):

F(s) = (3+s)/(s+1)(s^2+4s+13), then right?

OpenStudy (shamil98):

hold on, refreshing my memory on this stuff, haven't done these in a while..

OpenStudy (shamil98):

nvm, I can't remember what to do with the constant on the other side.. hold on, i'll call some higher-ups

OpenStudy (shamil98):

term* not constant

OpenStudy (yttrium):

Sure sure. Just wanna know, where i am getting wrong.

OpenStudy (shamil98):

y = e^(alpha*x) (c_1 cos Beta*x + c_2 sin Beta*x) that's the formula that i remember when dealing with a complex root. where alpha = -b/(2a) and beta = sqrt( 4ac-b^2)/(2a)

OpenStudy (shamil98):

all I got so far for the auxillary equation is: r^2 + 4r + 13 = 2e^(-t) what i'm not sure is that , does the 2e^(-t) become zero?..

OpenStudy (yttrium):

Ooops. I think were having different approach. I'm using laplace transformations here. No, 2e^(-t) = 2/(s+1) --> thru laplace transformation

OpenStudy (shamil98):

Oh, I'm using the formula for a second-order linear equation.. no wonder, i haven't learned that yet xD

OpenStudy (yttrium):

who can help me tho? T.T

zepdrix (zepdrix):

So you chose to solve this using Laplace Transform?

zepdrix (zepdrix):

sec lemme read the previous posts XD

OpenStudy (yttrium):

That's what I know, so far. But you still can teach me the other way.

zepdrix (zepdrix):

Mmm ya the Method of Undetermined Coefficients would be kind of nice for this one. You haven't learned about that yet? :o

OpenStudy (yttrium):

Ahhhhh. Already know that. I think that will be lot easier, right? But we are required to solve this using laplace transforms

zepdrix (zepdrix):

Oh ok ok fair enough. I'm a little rusty on taking inverse Laplace, but we'll see what we can do.

OpenStudy (yttrium):

See my above post. :)

zepdrix (zepdrix):

This step looks good so far. (s^2 + 4s + 13) F(s) = 3+s/(s+1) And this, ok ok good. F(s) = (3+s)/[(s+1)(s^2+4s+13)] So now we need to try and uhhhhhh... Mmm.. complete the square on the s maybe? And then Partial Fraction Decomposition afterwards.

OpenStudy (yttrium):

That's somewhat confuse me. Then it should be F(s) = (3+s)/[(s+1)((s+2)^2+9)], right?

zepdrix (zepdrix):

Mmmmmmmm ya that looks right. I'm trying to think if that's the approach we want.

zepdrix (zepdrix):

Ya I think we're on the right track. Remember how to do Partial Fraction Decomp?

OpenStudy (yttrium):

I've done that part, but I think it's getting messy. Well let's try again. (3+s)/[(s+1)(s^2+4s+13)] = A/(s+1) + (Cs+D)/(s^2+4s+13)

zepdrix (zepdrix):

You didn't use a B, which is kinda funny lol But setup looks good so far.

zepdrix (zepdrix):

Umm I would always try to stick in easy numbers before expanding it all out like that D: Remember how one of our terms had an (s+1) ? We can substitute s=-1 to easily solve for A. Hmm you should be getting 1/5's I think.. Not 1/4... Ahhhhhhh

zepdrix (zepdrix):

http://www.twiddla.com/1474743

zepdrix (zepdrix):

Darn I labeled them A B and C -_- Hopefully that's not confusing.

OpenStudy (yttrium):

What's B, then?

OpenStudy (yttrium):

B = -1/5 right?

zepdrix (zepdrix):

Ya I think so.

OpenStudy (yttrium):

You sure with that?

OpenStudy (yttrium):

But there's no wrong with my solution either.

zepdrix (zepdrix):

There probably is.. I don't have the attention span this late to check it though :'c I was able to check the expansion on wolfram, they seem to be coming up with the 1/5's also.

zepdrix (zepdrix):

`s: 1 = 4A +D` I think your s's should have given you: 1 = 4A + C + D

zepdrix (zepdrix):

3+s = A (s^2+4s+13) + Cs^2 + `C` + Ds +D 3+s = A (s^2+4s+13) + Cs^2 + `Cs` + Ds +D

OpenStudy (yttrium):

Yeah The hell

OpenStudy (yttrium):

3+s = A (s^2+4s+13) + Cs^2 + Cs + Ds +D Comparing coefficients: s^2: 0 = A+C A = -C s: 1 = 4A +C + D k: 3 = 13A + D eqn 2 to 3, sub eqn 1 to 3 1 = -3C + D 3 = -13C +D -2 = 10C C = -(1/5) A = 1/5 D = 2/5

OpenStudy (yttrium):

Is that right now?

OpenStudy (anonymous):

A Mathematica solution is attached.

zepdrix (zepdrix):

It seems like you're on the right track. I'm thinking the coefficient on sine might be off... Mmmmmm

OpenStudy (yttrium):

Then it will now be: f(t) = 1/5 [ 1/s+1 - s-2/[(s+2)^+9] f(t) = 1/15 [3e^(-t) - e^(-2t) (3cos3t - 2sin3t)] This is my final answer.

OpenStudy (yttrium):

Sorry but i don't have papers here. I'm just doing mental math :/

zepdrix (zepdrix):

Ok fair enough XD Job well done!

OpenStudy (yttrium):

Do we have same answers?